Would changing the volume of the original stock solution and the incremental dilution volumes to a new constant value affect the successive concentrations? (For example, starting with 10 mL and transferring in 10 mL increments?)
Answer #1

Suppose your stock solution has a concentration y M. If you take 10 mL of that solution to make a diluted solution then, the diluted solution contains, 10 mL of y M = 10y mLxM = 10y mLx(mol/L) = 10y (mL/L)xmol = 10y x 10-3 mol = 10y mmol of the substance with which you prepared the stock solution.

Now, say you are diluting to 100 mL. So, new concentration becomes (10y / 100) M = 0.1y M. For a successive same extent of dilution will yield 0.01y M solution.

Since, the concentration of a species is a function of the moles present in the solution, concentration has to be change with extent of dilution. For example, if we use 20 mL instead of 10 mL, then, concentration of the first solution becomes (20y / 100) M = 0.2y M and that of second solution becomes (20 x 0.2y / 100 ) M = 0.04y M. Indeed the concentration changed, because the no of moles are getting changed when we take different volumes of increments.